OpenStudy (anonymous):

I will give a medal follow and give medal to anyone who can help solve: log(5x+4)=2

OpenStudy (michele_laino):

what is the base of your logarithm?

OpenStudy (anonymous):

if it doesnt have a base, wouldn't the base be 10?

OpenStudy (michele_laino):

ok! If the base is 10, then, I apply the definition of logarithm and I can rewrite your equation as below: \[\begin{gathered} 5x + 4 = {10^2} \hfill \\ 5x = 100 - 4 \hfill \\ 5x = 96 \hfill \\ \end{gathered} \] what is x?

OpenStudy (anonymous):

19.2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

If the base is 2, then, again applying the definition of logarithm, we have: \[\begin{gathered} 5x + 4 = {2^2} \hfill \\ 5x = 4 - 4 \hfill \\ 5x = 0 \hfill \\ x = 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

but if there is no base shown, would the base automatically be 10?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

Ok thank you :)

OpenStudy (michele_laino):

thank you! :)